# RL-100: Performant Robotic Manipulation with Real-World Reinforcement Learning

### v1

@@ Abstract
Real-world robotic manipulation in homes and factories demands reliability, efficiency, and robustness that approach or surpass skilled human operators. We present RL-100, a real-world reinforcement learning training framework built on diffusion visuomotor policies trained bu supervised learning. RL-100 introduces a three-stage pipeline. First, imitation learning leverages human priors. Second, iterative offline reinforcement learning uses an Offline Policy Evaluation procedure, abbreviated OPE, to gate PPO-style updates that are applied in the denoising process for conservative and reliable improvement. Third, online reinforcement learning eliminates residual failure modes. An additional lightweight consistency distillation head compresses the multi-step sampling process in diffusion into a single-step policy, enabling high-frequency control with an order-of-magnitude reduction in latency while preserving task performance. The framework is task-, embodiment-, and representation-agnostic and supports both 3D point clouds and 2D RGB inputs, a variety of robot platforms, and both single-step and action-chunk policies. We evaluate RL-100 on seven real-robot tasks spanning dynamic rigid-body control, such as Push-T and Agile Bowling, fluids and granular pouring, deformable cloth folding, precise dexterous unscrewing, and multi-stage orange juicing. RL-100 attains 100% success across evaluated trials for a total of 900 out of 900 episodes, including up to 250 out of 250 consecutive trials on one task. The method achieves near-human teleoperation or better time efficiency and demonstrates multi-hour robustness with uninterrupted operation lasting up to two hours. The resulting policies generalize zero-shot to novel dynamics with an average success of 92.5% and adapt in a few-shot fashion to significant task variations, reaching an average of 86.7% after one to three hours of additional training. These results suggest a practical path to deployment-ready robot learning by starting from human priors, aligning training objectives with human-grounded metrics, and reliably extending performance beyond human demonstrations. For more results and videos, please visit our project website: https://lei-kun.github.io/RL-100/.
@@ 1 Introduction
Figure 1: Teaser. Real-robot snapshots illustrating the diversity of our task suite. Panels are ordered top-left $\rightarrow$ bottom-right: (a) Dynamic Push-T; (b) Agile Bowling; (c) Pouring. (d) Dynamic Unscrewing with a dexterous hand; (e) Dual-arm Folding; (f) Juicingstage 1 and stage 2 are shown side-by-side in the same panel; The grid highlights six of the seven tasks; juicing stage 1/2 are grouped into one panel for space.
Dexterous robotic manipulation stands as an iconic challenge in robotics (Cui and Trinkle, 2021; Luo et al., 2025). Real-world deployment beyond laboratories requires human-level reliability, efficiency, and robustness. Recent learning-based advances across generative diffusion policies (Chi et al., 2023; Ze et al., 2024), diffusion-based robot foundation models (Black et al., 2024a; Intelligence et al., 2025), sim-to-real RL (Yuan et al., 2025; Lin et al., 2025), and real-world RL (Luo et al., 2025, 2024a) have narrowed this gap, demonstrating human-like manipulation proficiency. In particular, generative policies and foundation models are trained or fine-tuned on high-quality, human-collected real-robot datasets at varying scales, providing strong human priors and enabling robots to acquire the efficient strategies used by skilled tele-operators. However, high-quality real-robot data remain scarce: teleoperation incurs perceptual and control latency and favors slow, conservative motions (Guo et al., 2025). Moreover, large-scale collection depends on skilled operators and is labor-intensive and expensive. As a result, state–action coverage is limited, undermining generalization and reliability. Consequently, this supervised paradigm is constrained by an imitation ceiling: under purely supervised objectives, performance is effectively bounded by demonstrator skill and inherits human inefficiencies, biases, and occasional errors.
@@
During policy rollouts, a human operator gives sparse success signals when needed, and the controller follows conservative operating limits. We use a unified policy-gradient objective across both iterative offline and online phases to fine-tune the diffusion sampler’s short-horizon denoising schedule (Song et al., 2021). This alignment yields stable updates across phases and strong fine-tuning sample efficiency. In addition, we interleave a lightweight distillation loss that compresses the $K$-step diffusion policy into a one-step consistency (Song et al., 2023) policy for deployment, reducing inference latency while maintaining or improving efficiency and robustness.
@@
    Deployment-centric results. On real robots, RL-100 achieves 100% reliability across seven tasks, including runs of up to 250 consecutive successes, attains efficiency comparable to or exceeding human teleoperation on multiple tasks, and demonstrates long-horizon robustness with uninterrupted operation for up to two hours, offering a promising route to deployment in homes and factories
@@ 2.1 Reinforcement Learning
Our post-training follows an offline-to-online paradigm. Following Lei et al. (2024), we employ a proximal policy optimization (PPO)-style objective (Schulman et al., 2017) to unify both stages. The core learning objective incorporates importance sampling for proximal policy updates:
@@ 2.2 Diffusion models
We overload the subscript $t$ from timestep in MDP to the step index in diffusion process in the following two subsections. Diffusion models (Ho et al., 2020a) learn to reverse a noising process that gradually corrupts clean data $x_{0}\!\in\!\mathbb{R}^{d}$ into Gaussian noise to reconstruct the original clean data distribution. Given a schedule $\{\alpha_{t}\}_{t=1}^{T}$ with $\alpha_{t}=1-\beta_{t}$, a sample $x_{0}$ drawn from the clean distribution, the forward noising process follows a closed form:
@@ 2.3 DDIM sampling with stochastic form
Denoising Diffusion Implicit Models (DDIM) (Song et al., 2021) provide a family of samplers that interpolate between deterministic and stochastic generation. Given a learned denoiser $\varepsilon_{\theta}$, the predicted clean sample at time $t$ is commonly written as
@@ 2.5 Diffusion policy and RL fine-tuning
Now we embed the $K$-step diffusion process as a sub-MDP into a single step in robotic manipulation MDP. Each denoising step can be viewed as sampling a cleaner noisy action $a^{\tau_{k-1}}$ from the Gaussian sub-policy in (7a). We therefore model the process as a $K$-step sub-MDP with:
@@ 3.2.1 Offline RL
by setting the updated policy as behavior policy: $\pi_{i+1}\!\coloneqq\!\pi$. Otherwise, we reject and keep the behavior policy unchanged ($\pi_{i+1}\!\coloneqq\!\pi_{i}$). In practice, we set $\delta=0.05\cdot|\widehat{J}^{\text{AM-Q}}(\pi_{i})|$ for adaptive thresholding. This OPE-gated rule yields conservative, monotonic behavior-policy improvement on $\mathcal{D}$.
@@ 4.1 Reinforcement Learning with Generative Diffusion Models
The integration of generative diffusion models with RL represents a paradigm shift in policy representation and optimization. Building upon foundational work in diffusion models (Ho et al., 2020b; Song et al., 2020) and flow matching (Lipman et al., 2023), recent advances have demonstrated the power of these generative frameworks in capturing complex, multimodal action distributions inherent in RL problems. Diffusion Q-Learning (DQL) (Wang et al., 2023b) pioneered this integration by replacing traditional Gaussian policies with conditional diffusion models in offline reinforcement learning, addressing fundamental limitations of parametric policies in modeling multimodal behaviors. This approach has evolved through multiple directions: weighted regression methods (Kang et al., 2023; Lu et al., 2023; Ding et al., 2024a) train diffusion policies through importance-weighted objectives to maximize learned Q-functions; reparameterization gradient approaches (Psenka et al., 2024; He et al., 2023; Ding et al., 2024b) utilize gradient-based optimization despite temporal backpropagation challenges; and sampling-based methods (Chen et al., 2023; Hansen-Estruch et al., 2024) provide effective but computationally expensive solutions. More recently, consistency-based extensions(Li et al., 2024) have further generalized diffusion and consistency policies to visual RL.
Recent works have also explored using RL to directly optimize diffusion models beyond traditional likelihood-based training. Black et al. (2023) demonstrate fine-tuning diffusion models with RL to maximize arbitrary reward functions, while Fan et al. (2024) apply similar techniques specifically to text-to-image generation using human feedback. Ren et al. (2024) introduce policy gradient methods tailored for diffusion-based policies, enabling effective online optimization while maintaining the expressiveness benefits of diffusion models.
@@ 4.2 Generative Diffusion Models in Robotics
The application of diffusion models to robotics has yielded transformative advances in visuomotor control, trajectory planning, and real-world deployment. Diffusion Policy (Chi et al., 2023) demonstrated breakthrough performance in visuomotor control by modeling action distributions as conditional diffusion processes, excelling at handling multimodal behaviors through visual conditioning and receding horizon control. This approach has inspired extensions including DiffClone (Sabatelli et al., 2024) and Diffusion Model-Augmented Behavioral Cloning (Wang et al., 2023a), which further improve upon traditional behavioral cloning by leveraging the expressiveness of diffusion models. The integration with pretrained visual representations, including R3M (Nair et al., 2022), VC-1 (Majumdar et al., 2023), and MoCo (He et al., 2020), has proven particularly effective for generalization across visual variations. Recent work on 3D Diffusion Policy (Ze et al., 2024) extends these capabilities to point cloud observations, while FlowPolicy (Chen et al., 2024) introduces consistency flow matching for robust 3D manipulation tasks. Lu et al. (2025) propose a triply-hierarchical diffusion policy that decomposes complex visuomotor tasks into multiple levels of abstraction, improving both learning efficiency and generalization.
Beyond direct policy learning, diffusion models have revolutionized trajectory planning in robotics. Janner et al. (2022) and Ajay et al. (2023) reformulate planning as conditional generation, producing complete state-action trajectories conditioned on rewards and constraints, excelling in long-horizon tasks requiring complex coordination. Shang et al. (2024) introduces an alternative paradigm by generating policy parameters rather than trajectories in latent spaces, offering computational advantages. Real-world deployment challenges have been addressed through methods like One-Step Diffusion Policy (Wang et al., 2024), which uses distillation to achieve real-time performance suitable for robotic control. Successful applications now span deformable object manipulation with PinchBot (Liu et al., 2024), multi-task learning, and navigation scenarios. The availability of large-scale datasets like BridgeData V2 (Walke et al., 2023) and foundation models such as $\pi_{0}$ (Black et al., 2024b) enables broader generalization across robotic platforms and tasks, accelerating the transition from laboratory demonstrations to practical robotic systems.
@@ 4.4 Real-world RL
These scattered advances converge in SERL (Luo et al., 2024b), a comprehensive framework that integrates high update-to-data ratio off-policy learning, automated resets, and visual reward specification to achieve several manipulation tasks. However, SERL relies solely on demonstrations and struggles with tasks requiring precision or recovery from failures. HIL-SERL (Luo et al., 2024c) addresses these limitations by incorporating real-time human corrections during training, enabling the policy to learn from mistakes and achieve perfect success rates across diverse tasks, including dual-arm coordination and dynamic manipulation.
While SERL and HIL-SERL report impressive on-robot learning efficiency and reliability on well-scoped tabletop skills, their evaluations typically employ action–space shaping (e.g., limiting wrist rotations and encouraging near-planar end-effector motion) and focus on short-horizon regimes with relatively low-dimensional control (Luo et al., 2024b, c). Such constraints are pragmatic for safety and sample efficiency, but they reduce policy expressivity and can cap performance on orientation-critical, contact-rich, or compositionally complex tasks. In everyday home and factory scenarios, many skills inherently require full SE(3) control and substantial reorientation, including: deformable manipulation with twist and regrasp (e.g., towel folding), insertion/ejection in confined cavities with large tilt changes (e.g., orange juicing), fluid and granular control that hinges on container tilt (e.g., controlled pouring), dynamic release and trajectory shaping (e.g., agile bowling), cable routing or cloth placement with out-of-plane rotations, and bimanual reorientation. [^1] We do not claim these systems cannot be extended to such settings; rather, the reported experiments emphasize rapid learning on well-scoped tasks, leaving broader generalization, long-horizon composition, and orientation-sensitive control less explored. In contrast, our system retains full 6-DoF control without hard rotation constraints and targets these under-explored regimes by (i) using a diffusion/consistency visuomotor policy to capture diverse human strategies, (ii) unifying offline-to-online improvement with an OPE-gated PPO-style objective for nearly-monotonic improvement, and (iii) enabling high-frequency control via one-step consistency distillation, across dual-arm, deformable, and dynamic tasks with larger cross-object generalization.
@@ 5.1 Overview
Key specifications for observation and action spaces of all tasks are summarized in Tab. 2. We randomize initial object placements for each trial to encourage policy generalization, with specific ranges for each task depicted in Fig. 5. Two types of reward structure are used for our tasks. For five tasks except Dynamic Push-T, we use a sparse reward function where the agent receives a reward of +1 upon successful task completion and 0 at all other timesteps, labeled by a human supervisor pressing a keyboard. The Dynamic Push-T task, which requires continuous, high-precision control, utilizes a dense shaped reward. The total reward at each timestep $t$ is $r_{t}=r_{\text{pose}}+r_{\text{static}}+r_{\text{smooth}}$, where $r_{\text{pose}}=\exp(-3e)-1$ punishes the SE(3) discrepancy $e$ between the T-block and the desired pose; $r_{\text{static}}=-1$ if the movement of T-block at one timestep is below a given threshold and otherwise 0; $r_{\text{smooth}}=-5\|a_{t}-a_{t-1}\|^{2}_{2}$ punishes jerky actions. More details about our setup (calibration, point-cloud processing, low-level control) are provided in supplementary materials.

### v2

@@ Abstract
Real-world robotic manipulation in homes and factories demands reliability, efficiency, and robustness that approach or surpass skilled human operators. We present RL-100, a real-world reinforcement learning training framework built on diffusion visuomotor policies trained by supervised learning. RL-100 introduces a three-stage pipeline. First, imitation learning leverages human priors. Second, iterative offline reinforcement learning uses an Offline Policy Evaluation procedure, abbreviated OPE, to gate PPO-style updates that are applied in the denoising process for conservative and reliable improvement. Third, online reinforcement learning eliminates residual failure modes. An additional lightweight consistency distillation head compresses the multi-step sampling process in diffusion into a single-step policy, enabling high-frequency control with an order-of-magnitude reduction in latency while preserving task performance. The framework is task-, embodiment-, and representation-agnostic and supports both 3D point clouds and 2D RGB inputs, a variety of robot platforms, and both single-step and action-chunk policies. We evaluate RL-100 on seven real-robot tasks spanning dynamic rigid-body control, such as Push-T and Agile Bowling, fluids and granular pouring, deformable cloth folding, precise dexterous unscrewing, and multi-stage orange juicing. RL-100 attains 100% success across evaluated trials for a total of 900 out of 900 episodes, including up to 250 out of 250 consecutive trials on one task. The method achieves near-human teleoperation or better time efficiency and demonstrates multi-hour robustness with uninterrupted operation lasting up to two hours. Policies generalize zero-shot to novel dynamics with an average 90.0% success, adapt in a few-shot manner to significant task variations (86.7% after 1–3 hours of additional training), and remain robust under aggressive human perturbations (about 95% success). Notably, our juicing robot served random customers continuously for about 7 hours without failure when zero-shot deployed in a shopping mall. These results suggest a practical path to deployment-ready robot learning by starting from human priors, aligning training objectives with human-grounded metrics, and reliably extending performance beyond human demonstrations. For more results and videos, please visit our project website: https://lei-kun.github.io/RL-100/.
@@ 1 Introduction
Figure 1: Teaser. Real-robot snapshots illustrating the diversity of our task suite. Panels are ordered top-left $\rightarrow$ bottom-right: (a) Dynamic Push-T; (b) Agile Bowling; (c) Pouring. (d) Dynamic Unscrewing with a dexterous hand; (e) Dual-arm Folding; (f) Juicingstage 1 and stage 2 are shown side-by-side in the same panel; The grid highlights all seven tasks; juicing stage 1/2 are grouped into one panel for space.
Dexterous robotic manipulation stands as an iconic challenge in robotics (Cui and Trinkle, 2021; Luo et al., 2025). Real-world deployment beyond laboratories requires human-level reliability, efficiency, and robustness. Recent learning-based advances across generative diffusion policies (Chi et al., 2023; Ze et al., 2024), diffusion-based robot foundation models (Black et al., 2024a; Intelligence et al., 2025), sim-to-real RL (Yuan et al., 2025; Lin et al., 2025), and real-world RL (Luo et al., 2025, 2024) have narrowed this gap, demonstrating human-like manipulation proficiency. In particular, generative policies and foundation models are trained or fine-tuned on high-quality, human-collected real-robot datasets at varying scales, providing strong human priors and enabling robots to acquire the efficient strategies used by skilled tele-operators. However, high-quality real-robot data remain scarce: teleoperation incurs perceptual and control latency and favors slow, conservative motions (Guo et al., 2025). Moreover, large-scale collection depends on skilled operators and is labor-intensive and expensive. As a result, state–action coverage is limited, undermining generalization and reliability. Consequently, this supervised paradigm is constrained by an imitation ceiling: under purely supervised objectives, performance is effectively bounded by demonstrator skill and inherits human inefficiencies, biases, and occasional errors.
@@
During policy rollouts, a human operator gives sparse success signals when needed, and the controller follows conservative operating limits. We use a unified policy-gradient objective across both iterative offline and online phases to fine-tune the diffusion sampler’s short-horizon denoising schedule (Ho et al., 2020a). This alignment yields stable updates across phases and strong fine-tuning sample efficiency. In addition, we interleave a lightweight distillation loss that compresses the $K$-step diffusion policy into a one-step consistency (Song et al., 2023) policy for deployment, reducing inference latency while maintaining or improving efficiency and robustness.
@@
    Deployment-centric results. On real robots, RL-100 achieves 100% success across seven tasks (including runs of up to 250 consecutive trials), and matches or exceeds human teleoperation efficiency on multiple tasks. Notably, the juicing robot served random customers continuously for about 7 hours without failure under zero-shot shopping mall deployment.
@@ 2.1 Reinforcement Learning
Our post-training follows an offline-to-online paradigm. Following LEI et al. (2024), we employ a proximal policy optimization (PPO)-style objective (Schulman et al., 2017; Zhuang et al., 2023) to unify both stages. The core learning objective incorporates importance sampling for proximal policy updates:
@@ 2.2 Diffusion models
We overload the subscript $t$ from timestep in MDP to the step index in diffusion process in the following two subsections. Diffusion models (Ho et al., 2020b) learn to reverse a noising process that gradually corrupts clean data $x_{0}\!\in\!\mathbb{R}^{d}$ into Gaussian noise to reconstruct the original clean data distribution. Given a schedule $\{\alpha_{t}\}_{t=1}^{T}$ with $\alpha_{t}=1-\beta_{t}$, a sample $x_{0}$ drawn from the clean distribution, the forward noising process follows a closed form:
@@ 2.3 DDIM sampling with stochastic form
Denoising Diffusion Implicit Models (DDIM) (Ho et al., 2020a) provide a family of samplers that interpolate between deterministic and stochastic generation. Given a learned denoiser $\varepsilon_{\theta}$, the predicted clean sample at time $t$ is commonly written as
@@ 2.5 Diffusion policy and RL fine-tuning
Now we embed the $K$-step diffusion process as a sub-MDP into a single step in robotic manipulation MDP (Ren et al., 2024). Each denoising step can be viewed as sampling a cleaner noisy action $a^{\tau_{k-1}}$ from the Gaussian sub-policy in (7a). We therefore model the process as a $K$-step sub-MDP with:
@@ 3.2.1 Offline RL
by setting the updated policy as behavior policy: $\pi_{i+1}\!\coloneqq\!\pi$. Otherwise, we reject and keep the behavior policy unchanged ($\pi_{i+1}\!\coloneqq\!\pi_{i}$). In practice, we set $\delta=0.05\cdot|\widehat{J}^{\text{AM-Q}}(\pi_{i})|$ for adaptive thresholding. This OPE-gated rule yields conservative, monotonic behavior-policy improvement on $\mathcal{D}$. Given the estimated advantage, the loss function of offline RL $\mathcal{L}^{\text{off}}_{\text{RL}}$ equals to Eq. (19).
@@ 4.1 Reinforcement Learning with Generative Diffusion Models
The integration of generative diffusion models with RL represents a paradigm shift in policy representation and optimization. Building upon foundational work in diffusion models (Ho et al., 2020a; Song et al., 2020) and flow matching (Lipman et al., 2023), recent advances have demonstrated the power of these generative frameworks in capturing complex, multimodal action distributions inherent in RL problems. Diffusion Q-Learning (DQL) (Wang et al., 2023) pioneered this integration by replacing traditional Gaussian policies with conditional diffusion models in offline reinforcement learning, addressing fundamental limitations of parametric policies in modeling multimodal behaviors. This approach has evolved through multiple directions: weighted regression methods (Kang et al., 2023; Lu et al., 2023; Ding et al., 2024) train diffusion policies through importance-weighted objectives to maximize learned Q-functions; reparameterization gradient approaches (Psenka et al., 2024; He et al., 2023; Ding and Jin, 2023) utilize gradient-based optimization despite temporal backpropagation challenges; and sampling-based methods (Chen et al., 2024) provide effective but computationally expensive solutions. More recently, consistency-based extensions(Li et al., 2024) have further generalized diffusion and consistency policies to visual RL.
Recent works have also explored using RL to directly optimize diffusion models beyond traditional likelihood-based training. Black et al. (2024b) demonstrate fine-tuning diffusion models with RL to maximize arbitrary reward functions, while Fan et al. (2023) apply similar techniques specifically to text-to-image generation using human feedback. Ren et al. (2024) introduce policy gradient methods tailored for diffusion-based policies, enabling effective online optimization while maintaining the expressiveness benefits of diffusion models.
@@ 4.2 Generative Diffusion Models in Robotics
The application of diffusion models to robotics has yielded transformative advances in visuomotor control, trajectory planning, and real-world deployment. Diffusion Policy (Chi et al., 2023) demonstrated breakthrough performance in visuomotor control by modeling action distributions as conditional diffusion processes, excelling at handling multimodal behaviors through visual conditioning and receding horizon control. This approach has inspired extensions including DiffClone (Mani et al., 2024) and Diffusion Model-Augmented Behavioral Cloning (Chen et al., 2023), which further improve upon traditional behavioral cloning by leveraging the expressiveness of diffusion models. The integration with pretrained visual representations, including R3M (Nair et al., 2022), VC-1 (Majumdar et al., 2023), and MoCo (He et al., 2020), has proven particularly effective for generalization across visual variations. Recent work on 3D Diffusion Policy (Ze et al., 2024) extends these capabilities to point cloud observations, while FlowPolicy (Zhang et al., 2025a) introduces consistency flow matching for robust 3D manipulation tasks. Lu et al. (2025) propose a triply-hierarchical diffusion policy that decomposes complex visuomotor tasks into multiple levels of abstraction, improving both learning efficiency and generalization.
Beyond direct policy learning, diffusion models have revolutionized trajectory planning in robotics. Janner et al. (2022) and Ajay et al. (2022) reformulate planning as conditional generation, producing complete state-action trajectories conditioned on rewards and constraints, excelling in long-horizon tasks requiring complex coordination. Hegde et al. (2025) introduces an alternative paradigm by generating policy parameters rather than trajectories in latent spaces, offering computational advantages. Real-world deployment challenges have been addressed through methods like One-Step Diffusion Policy (Wang et al., 2025), which uses distillation to achieve real-time performance suitable for robotic control. Successful applications now span deformable object manipulation with PinchBot (Bartsch et al., 2025), multi-task learning, and navigation scenarios. The availability of large-scale datasets like BridgeData V2 (Walke et al., 2023) and foundation models such as $\pi_{0}$ (Black et al., 2024a) enables broader generalization across robotic platforms and tasks, accelerating the transition from laboratory demonstrations to practical robotic systems.
@@ 4.4 Real-world RL
These scattered advances converge in SERL (Luo et al., 2024), a comprehensive framework that integrates high update-to-data ratio off-policy learning, automated resets, and visual reward specification to achieve several manipulation tasks. However, SERL relies solely on demonstrations and struggles with tasks requiring precision or recovery from failures. HIL-SERL (Luo et al., 2025) addresses these limitations by incorporating real-time human corrections during training, enabling the policy to learn from mistakes and achieve perfect success rates across diverse tasks, including dual-arm coordination and dynamic manipulation.
While SERL and HIL-SERL report impressive on-robot learning efficiency and reliability on well-scoped tabletop skills, their evaluations typically employ action–space shaping (e.g., limiting wrist rotations and encouraging near-planar end-effector motion) and focus on short-horizon regimes with relatively low-dimensional control (Luo et al., 2024, 2025). Such constraints are pragmatic for safety and sample efficiency, but they reduce policy expressivity and can cap performance on orientation-critical, contact-rich, or compositionally complex tasks. In everyday home and factory scenarios, many skills inherently require full SE(3) control and substantial reorientation, including: deformable manipulation with twist and regrasp (e.g., towel folding), insertion/ejection in confined cavities with large tilt changes (e.g., orange juicing), fluid and granular control that hinges on container tilt (e.g., controlled pouring), dynamic release and trajectory shaping (e.g., agile bowling), cable routing or cloth placement with out-of-plane rotations, and bimanual reorientation. [^1] We do not claim these systems cannot be extended to such settings; rather, the reported experiments emphasize rapid learning on well-scoped tasks, leaving broader generalization, long-horizon composition, and orientation-sensitive control less explored. In contrast, our system retains full 6-DoF control without hard rotation constraints and targets these under-explored regimes by (i) using a diffusion/consistency visuomotor policy to capture diverse human strategies, (ii) unifying offline-to-online improvement with an OPE-gated PPO-style objective for nearly-monotonic improvement, and (iii) enabling high-frequency control via one-step consistency distillation, across dual-arm, deformable, and dynamic tasks with larger cross-object generalization.
@@ 5.1 Overview
Key specifications for observation and action spaces of all tasks are summarized in Tab. 2. We randomize initial object placements for each trial to encourage policy generalization, with specific ranges for each task depicted in Fig. 5. Two types of reward structure are used for our tasks. For all tasks except Dynamic Push-T, we use a sparse reward function where the agent receives a reward of +1 upon successful task completion and 0 at all other timesteps, labeled by a human supervisor pressing a keyboard. The Dynamic Push-T task, which requires continuous, high-precision control, utilizes a dense shaped reward. The total reward at each timestep $t$ is $r_{t}=r_{\text{pose}}+r_{\text{static}}+r_{\text{smooth}}$, where $r_{\text{pose}}=\exp(-3e)-1$ punishes the SE(3) discrepancy $e$ between the T-block and the desired pose; $r_{\text{static}}=-1$ if the movement of T-block at one timestep is below a given threshold and otherwise 0; $r_{\text{smooth}}=-5\|a_{t}-a_{t-1}\|^{2}_{2}$ punishes jerky actions. More details about our setup (calibration, point-cloud processing, low-level control) are provided in supplementary materials.